OpenStudy (lovelyharmonics):

Give an example of a rational function that has a horizontal asymptote of y = 2/9.

ganeshie8 (ganeshie8):

use below : \(\large \frac{\color{red}{a}x+b}{\color{red}{c}x+d}\) has a horizontal asymptote af \(y = \frac{\color{red}{a}}{\color{red}{c}}\)

OpenStudy (anonymous):

give the leading coefficient of the numerator a 2, and the leading coefficient of the denominator a 9 and write whatever you want top and bottom so long as the degrees are the same

OpenStudy (anonymous):

you could write what @ganeshie8 wrote, or you could write \[y=\frac{2x^2+3x-4}{9x^2-5x}\] or anything you like

OpenStudy (lovelyharmonics):

2x+b / 9x+d whats b and d?

OpenStudy (anonymous):

makes no difference all you care is that the degrees are the same and the ratio of the leading coefficients is \(\frac{2}{9}\)

ganeshie8 (ganeshie8):

just put some numbers for b and d

OpenStudy (lovelyharmonics):

okay and x?

OpenStudy (anonymous):

or be a wise guy and write \[\frac{2x^{37}+1}{9x^{37}-5}\]

OpenStudy (lovelyharmonics):

.-.what?

OpenStudy (ipwnbunnies):

Lol, good job satellite.

OpenStudy (anonymous):

make the degree of the top the same as the degree of the bottom and the ratio of the leading coefficients \(\frac{2}{9}\) @ganeshie8 had the degree of the top and bottom both 1, i.e. \(\frac{2x+something}{9x+whatever}\) but it makes no difference so long as the degrees are the same

OpenStudy (lovelyharmonics):

2x+1 / 9x+1?

OpenStudy (anonymous):

that will certainly work

OpenStudy (lovelyharmonics):

yay thanks ^.^

OpenStudy (anonymous):

yw